OpenStudy (anonymous):

Find lim x goes to 0 for x square over 2-square root of x+4 and lim x goes to 2 for xsquare -x-2 over x square -4 and lim goes to 0 for sin square x over xsquare

OpenStudy (dumbcow):

wow thats a mouthful, can you post each limit separately using equation editor? it help clarify for people helping you thx

OpenStudy (anonymous):

I'm posting this from my phone m. Sorry

OpenStudy (anonymous):

I really need help I only got two hours before mid trrm exam

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{x^2}{2 \sqrt{x+4}}\] like this?

OpenStudy (anonymous):

or subtract the square root?

OpenStudy (anonymous):

You subtract it

hartnn (hartnn):

\(\Large \lim \limits _{x \rightarrow 0} \frac{x^2}{2 -\sqrt{x+4}}\) do you know how to rationalize the denominator ?

OpenStudy (anonymous):

No

hartnn (hartnn):

multiply and divide by conjugate of denominator, that is \(2+\sqrt{x+4}\) what do u get ?

OpenStudy (anonymous):

The answer is 8 +4 square root of 6 over 6 is it correct ?

hartnn (hartnn):

no, thats not correct. how did you get that ?

OpenStudy (anonymous):

WE MILTIPLIED the conjugate .. I want the answers to anlazy it myself

hartnn (hartnn):

\(\large \lim \limits _{x \rightarrow 0} \dfrac{x^2}{2 -\sqrt{x+4}} \dfrac{2 +\sqrt{x+4}}{2 +\sqrt{x+4}} = \dfrac{(2 -\sqrt{x+4)}x^2}{4- (x+4)} \\ \large =\dfrac{x^\cancel{2}( 2 -\sqrt{x+4})}{-\cancel x} = \dfrac{x (2 -\sqrt{x+4})}{-1}\) now you can just plug in x =0 in this! what do u get ?

OpenStudy (anonymous):

Thanks man , you helped a lot ^^

hartnn (hartnn):

welcome ^_^